Meeting:2011/03/24

From SOWNWiki
Jump to: navigation, search

Meeting (to be) held on 2011/03/24 at 19:00 in Zepler CLS Lecture Room

Previous meeting: 17 March 2011 19:00:00 Next meeting: 31 March 2011 19:00:00


Minutes

  • Icinga Alerts (Crwilliams)
    • SUSU / MEM (Leth)
      • Is it possible to upgrade syslog-ng?
      • morse/Davetaz happy for Leth to do this
      • Actually don't upgrade because libc needs upgrading that will require all binaries to be updated which is really not a good idea.
    • SUSU / Disk /tmp/ (Leth)
      • commit-log has been cleared after copy made to auth
    • AUTH/AUTH
      • Not sure what is failing yet
    • AUTH/EAPOL-EDUROAM
      • We need to add checks over PEAP that will be provided by mike
      • Need to investigate how many protocols are actually supported by eduroam
    • All Homenode tunnels outage - OpenVPN (SSL) (Davetaz/daveruss)
      • Updating openvpn package broke all the tunnels
      • Problem caused by temporary client config files not been able to write to disk because of permissions
      • We were using a debug flag that we shouldn't be using
      • Reduced run level of OpenVPN from 3 to 2
      • Specified temp directory as /tmp so that openvpn can write to disk as nobody
      • Andrew_W said we need to update our wiki pages on OpenVPN top reflect these customisations
    • MONITOR/DISK / - Database issues?
      • Leth setup manual logrotate of syslog-ng
    • AUTH/DISK / - Database issues?
      • Davetaz moved all the mysql binary logs to the /home partition in /home/large_directories/
      • We need to change the config for mysql so that it doesn't make binary logs as we don't need them
      • Look at the mysql logging and see if we can find a way of tidying this up on the nightly backups
    • GW/DISK /
      • Leth probably gzipped some stuff
      • Some log files could probably be cleared out
      • Nagios was running on GW, just logging itself without telling anyone. This has been stopped but needs to be removed and its log files cleared up
    • 32/3089-UPS - remove from Icinga?
      • Davetaz happy for this to be removed
      • Reporting will be made directly to email and irc
      • crwilliams reckons we should log the data in icinga but not monitor or warn
    • MONITOR/DEBSUMS
      • Timing out so timeout increased.
      • Taking too long, needs to be investigated.
    • MONITOR/check_nsca* (Leth)
      • Has been flapping. Leth has now fixed both of these.
    • PACKAGE warnings not reported in the channel
      • crwilliams is not sure why this is happening.


  • Cron emails (Leth)
    • Forwarding - Not enabled on GW. Other servers?
      • Davetaz will add the relay config for mail on the other servers as he has got the config sitting around
    • ntp
      • Fixed by Leth and morse
    • logrotate
      • Needs to be looked into
    • lost+found
      • Full of files that need to be investigated/deleted to stop email being sent out


  • data.southampton.ac.uk
    • SOWN access points info is now published here.
    • Further work needs to be done to represent this information semantically.


  • Admin Tasks
    • Undercroft situation (Crwilliams)
      • No progress
    • Kit order (Davetaz)
      • Backup server disks
        • Received present from AndyLandy
    • DNS records (Davetaz)
      • Need to chase AndyLandy about this
    • Updates to mailing lists/aliases (daveruss)
      • lists updated
    • SUWS webspace at susu.org (Scorpia)
      • suws.susu.org requested
      • Might be worth asking for a dns pointer to something in ECS (probably websites1)
    • Constitution changes (Scorpia)
      • Scorpia has to attend a meeting with SUSU about these changes on 3rd May 2011



  • Ham Radio (HeliosFA)
    • OfCom replied to email but didn't really answer the question
    • Want to get old call signs back but can't track down old registrants
    • Recovered linear amplifier from Dennis
    • No risk assessment yet for going on Zepler roof for investigation
    • Mike has old bank statements for SOWN/SUWS


  • Workshop Review
    • Wifi Survey (daveruss)
      • Two machines have the scanning software on
      • Need to get UGs to take them and run tests + see if there are any bugs with the app
    • Icinga Deployment (Crwilliams)
      • Didn't bother copying log file database tables as we have the log files
      • Deployment is now complete and Nagios can now be turned off
      • Will do that next week at start of meeting
    • Eduroam Campus Wide (Davetaz)
      • Mike ran some tests from the eduroam website. Didn't see any packets going on/off radius server whilst test running
      • daveruss couldn't get eduroam working from Oxford connecting to soton.ac.uk and test.soton.ac.uk (and ecs.soton.ac.uk)
    • Inventory
      • Nothing major found to be missing or broken except Yagi that has been missing a long time.
      • Added boxes to database so we can refer to location. This needs to be linked up a bit better.




  • AOB
    • /opt/sown/{bin,sbin}
      • Not discussed
Facts about "2011/03/24"
Has date24 March 2011 19:00:00 +
Has end date24 March 2011 20:00:00 +
Has location59/1257 +